# CReFT-CAD: Boosting Orthographic Projection Reasoning for CAD via Reinforcement Fine-Tuning

### v1

@@ 4.1.1 Task 1: Dichotomous choice task
As for this task, the model’s output is constrained to dichotomous responses, either "yes" or "no." A "yes" answer is given only when the entire 15-dimensional parameter space is completely correct, while any error, even in a single parameter, results in a "no" response. This strict criterion ensures that the model is incentivized to produce fully accurate parameterizations. To align with this setup, the reward function is defined as follows:
@@ 4.1.2 Task 2: Multiple choice task
Training strategy. In the task 2, we introduce multiple choice tasks to further increase the complexity of decision-making. These tasks require the model to select from multiple parameter value lists, which significantly enhancing the cognitive load compared to binary decisions. Building on the first stage, we pre-sample a subset of image-text instruction pairs to perform instruction-following training on the LLM. This training allows the model to generate structured outputs in the form of multiple choice questions, enabling the reinforcement learning reward mechanism. Additionally, we designed a stringent reward function that encourages the model to maximize the utilization of image-text instruction pairs, ensuring that it extracts and applies all relevant information from both the visual and textual components. The reward function incentivizes the model to select the most accurate option from the provided set of choices. Any incorrect selection results in a reward of 0, thereby pushing the model to focus on precision and reliability in its decision-making process.
@@
$R_{\text{P2}}=\begin{cases}1&\text{if }S_{\text{selected}}=S_{\text{correct}}\\ 0.2&\text{if }S_{\text{selected}}\supseteq S_{\text{correct}}\text{ but }S_{\text{selected}}\neq S_{\text{correct}}\\ 0&\text{if }(S_{\text{selected}}\cap S_{\text{correct}}\neq\emptyset)\text{ and }(S_{\text{selected}}\cap S_{\text{incorrect}}\neq\emptyset)\\ 0&\text{if }S_{\text{selected}}=S_{\text{incorrect}}\end{cases}$ (2)
Where $S_{\text{Selected}}=S_{\text{correct}}$ means the model selects only the correct parameter value lists; $S_{\text{Selected}}\supseteq S_{\text{correct}}\text{ but }S_{\text{Selected}}\neq S_{\text{correct}}$ means the model selects all correct lists but fails to select all of them; $S_{\text{Selected}}\cap S_{\text{correct}}\neq\emptyset$ and $S_{\text{Selected}}\cap S_{\text{incorrect}}\neq\emptyset$ means the model selects both correct and incorrect parameter value lists; $S_{\text{Selected}}=S_{\text{incorrect}}$ means the model selects incorrect parameter value lists.
@@ 4.1.3 Task 3: Parameterization tasks based on chain-of-thought
Training strategy. In the training process, we begin with the model trained in task 2, and sample synthetic data to construct standard orthographic projection reasoning question-answer pairs. These pairs are then used to perform small-batch fine-tuning. Once the model is fine-tuned, we assess the difficulty of predicting the 15-dimensional parameters by evaluating the model’s performance on a test set. The prediction accuracy for each parameter is categorized into three difficulty levels: Easy Attributes: Attributes with an accuracy greater than 0.8. Medium Attributes: Attributes with accuracy between 0.2 and 0.8. Difficult Attributes: Attributes with accuracy less than 0.2. This difficulty classification is then used to design the reward function The reward for each correctly answered attribute is determined by its difficulty level as follows:
$R_{\text{P3}}=\begin{cases}1&\text{if the model correctly predicts an easy attribute }\\ 1.5&\text{if the model correctly predicts a medium attribute }\\ 2&\text{if the model correctly predicts a hard attribute }\\ 0&\text{if the model's prediction is incorrect }\end{cases}$ (3)
@@ 4.2 Supervised post-tuning
After the CReFT stage, where the model developed robust reasoning and generalization capabilities, the focus of the Supervised Post-tuning task shifts to enhancing the model’s instruction-following ability. During this stage, we rethink the industrial design application context and, with a user-centric approach, reformulate tasks in the style of Visual Question Answering (VQA). By leveraging multimodal language models, we design specific tasks that guide the model to output a comprehensive list of parameters and make critical judgments based on orthographic projection. In this stage, the model is trained to process the following tasks: 1) Full Parameter List Output Task: The goal is to generate a fully accurate set of parameter key-value pairs, which is the most common format for outputs during the simulation phase. 2) Parameter Validation Task: The task is to compare the extracted parameter key-value pairs with the accurate parameter key-value pairs in the orthographic projection to determine whether the extracted parameters are correct. 3) View Matching Task: The model is given two individual orthographic projection views. The model must, based on parameter constraints and rasterized images, assess whether the views correspond to the same overall 3D object or design. 4) Component Counting Task: The model is required to identify and count specific components in the drawing, such as bridge piers or columns, based on the given parameter constraints. The output of this task helps to quantify the number of elements in the design, aiding in material estimation and component verification.
By employing this post-tuning phase, we enhance the model’s ability to interpret and respond to complex design queries, thereby improving its interactive reasoning abilities and making it more applicable to real-world industrial design workflows.
@@ 5 Experiments
Implementation Details. All experiments are conducted on NVIDIA A100 GPUs. Most experiments use GOT-OCR2.0 as the base model, trained on a single server equipped with 8 A100 GPUs and a batch size of 64. We use the AdamW optimizer with a cosine annealing scheduler. The hyperparameters are configured as follows: (1) Learning rate: 1e-6 for RL (GRPO) training and 2e-5 for baseline SFT experiments. (2) Maximum input image size: 2,483,776 pixels. (3) Total GRPO training steps: 1500.
Datasets and Metrics. We evaluate both the leading vision–language models (VLMs) and the proposed method on TriView2CAD, following the evaluation strategy outlined in Sec. 3.2. The synthetic test set is used to assess in-domain accuracy. Due to the difficulty in obtaining real-world data, all 3,000 real-world samples are reserved exclusively for testing, allowing us to evaluate out-of-distribution (OOD) generalization in real-world scenarios.
We design four distinct prompt configurations to probe their orthographic projection reasoning capabilities: Test image-Only: Models receive only the target rasterized drawings. Test image + Reference Image: In addition to the target drawings, models are given a reference image—rendered on the same geometry layer—with all dimension names to supply semantic alignment. Test image + Answered Pair: Inputs consist of the target drawings plus a correctly paired raster image and its parameter vector, providing an exemplar mapping between visual features and quantitative parameters. Test image + Atttribute Explanation: Models are supplied with the target drawings alongside a detailed textual interpretation authored by professional engineers.
Given that orthographic projection reasoning involves complex multi-step inference and current pretrained VLMs limit exposure to CAD drawings, we incorporate explicit reasoning guidance into each prompt format to encourage clear reasoning trajectories. For each of the four different prompt input formats, we outline detailed step-by-step instructions for completing the task. These steps are designed to guide the model through the reasoning process in a structured manner. The complete set of instructions for each format is provided in the supplementary materials.
@@
5.1 Performance on in-domain test set.
The experimental results, as shown in Tab. 3, demonstrate that our method significantly outperforms existing VLMs on orthographic projection reasoning tasks across all prompt formats. Specifically, in the Without Reasoning Guidance setting, the "+Answered Pair" prompt achieves the highest accuracy of 83.24%. In the Reasoning Guidance setting, our model achieves an accuracy of 84.03% in the "+Attribute Explanation" prompt. Furthermore, our method exhibits strong performance even in the "Test Image Only" condition, where we achieve 80.86%. This improvement highlights the effectiveness of our approach.
@@
5.2 Performance on real-world test set.
We utilize all 3,000 real-world samples to evaluate out-of-distribution (OOD) generalization in real-world scenarios. The baseline model, Qwen2.5-VL, achieves an accuracy of 13.47%, which is lower than its performance on synthetic data (24.54%). This performance drop is primarily due to the more complex structure and occlusion relationships present in the real-world data, as discussed in Sec. 3.1, which increases the difficulty of the task. We also compare the results with a model trained using a SFT strategy, which achieve an accuracy of 36.15%. This improvement can be attributed to significant gains in Primitive Counting tasks. However, the model still struggle with Composite Parameter Computation, highlighting the limitations of traditional SFT. Our method, using CReFT, achieves an accuracy of 46.67%, demonstrating that our method can effectively address OOD generalization and lead to significant improvements on more complex real-world data.
@@ 5.3 Abalation study
With CoT VS without CoT. The results from Tab. 2 reveal a significant improvement in performance when incorporating chain-of-thought (CoT) reasoning. Specifically, the accuracy for individual Task 3 increase from 46.16% to 74.15% after adding CoT , demonstrating a clear performance boost. When combining multiple training tasks, such as Task 1 + Task 3 and Task 2 + Task 3, the improvements are even more pronounced, with accuracies reaching 77.90% and 67.75%, respectively. The most notable enhancement is observed in the combination of Task 1 + Task 2 + Task 3, where the accuracy increases from 46.24% to 81.35%. The primary source of these improvements lies in how CoT helps break down complex Composite Parameter Computation tasks into more manageable, incremental steps. The CoT mechanism, however, guides the model through each individual component of the computation, allowing for a step-by-step approach that improves the model’s reasoning capability.
@@ Appendix B Further analysis regarding the three findings
1) Remains a Tough Challenge for pretrained VLMs. Tab. 3 reports performance across four prompt formats without and with reasoning guidance. Overall, orthographic projection reasoning entails not only reading textual dimensions and visual features, but also matching annotations to geometry primitives, counting structural elements, and computing composite parameters. Consequently, none of the off-the-shelf models fully masters this composite task. Qwen2.5-VL [42] achieves the highest accuracy—38.88% under reasoning guidance with the Test Image + Attribute Explanation prompt. This improvement stems primarily from its superior ability to parse geometry layers and read dimension labels. Crucially, its performance on reasoning-intensive parameters remains low. These findings underscore that orthographic projection CAD cannot be solved by prompting alone and requires dedicated fine-tuning strategy.
@@
3)Introducing reasoning guidance yields consistent gains across all seven VLMs. Introducing reasoning guidance demonstrates that step-wise reasoning guidance helps models better decompose the multi-step orthographic projection tasks. In the absence of reasoning guidance, visual exemplar prompts (+Reference Image and +Answered Pair) deliver the largest relative improvements. However, when reasoning guidance is added, the Attribute Explanation prompt shows the greatest uplift. It stems from reasoning guidance is textual form, which synergizes most effectively with textual inputs. Hence, models with stronger text encoders (e.g., DeepSeek-V1) exhibit disproportionately larger boosts. These results underscore the necessity of integrating structured reasoning guidance to advance orthographic projection CAD reasoning.

### v2

@@ 4.1.1 Task 1: Dichotomous choice task
For this task, the model’s output is restricted to dichotomous responses: "yes" or "no." A "yes" response is returned if and only if all parameters in the 15-dimensional parameter space are fully accurate, whereas any error—even in a single parameter—triggers a "no" response. This stringent criterion incentivizes the model to generate fully accurate parameterizations. Consistent with this design, the reward function is defined as follows:
@@ 4.1.2 Task 2: Multiple choice task
Training strategy. Task 2 introduces multiple-choice tasks to further increase decision-making complexity. These tasks require the model to select from multiple parameter value lists, which significantly enhances the cognitive load compared to binary decisions. Building on the first task, we pre-sample a subset of image-text instruction pairs to conduct instruction-following training on the LLM. This training enables the model to generate structured outputs in the form of multiple-choice questions, facilitating the activation of the reinforcement learning (RL) reward mechanism. Additionally, we design a stringent reward function that encourages the model to fully utilize image-text instruction pairs, ensuring it extracts and leverages all relevant information from both visual and textual modalities. The reward function incentivizes the model to select the most accurate option from the provided choices. Any incorrect selection yields a reward of 0, compelling the model to prioritize precision and reliability in its decision-making.
@@
[…]
@@ 4.1.3 Task 3: Parameterization tasks based on chain-of-thought
Training strategy. For this task’s training, we start with the model fine-tuned in Task 2 and sample synthetic data to construct standard orthographic projection reasoning QA pairs. These pairs are used for small-batch fine-tuning. After fine-tuning, we assess the difficulty of predicting the 15-dimensional parameters by evaluating the model’s performance on a held-out test set. Each parameter’s prediction accuracy is categorized into three difficulty levels: Easy Attributes: Attributes with an accuracy greater than 0.8. Medium Attributes: Attributes with accuracy between 0.2 and 0.8. Difficult Attributes: Attributes with accuracy less than 0.2. This difficulty classification is subsequently used to design the reward function. The reward for each correctly predicted attribute is determined by its difficulty level as follows:
[…]
@@ 4.2 Supervised post-tuning
Following the CReFT stage—during which the model has developed robust reasoning and generalization capabilities—the focus of Supervised Post-tuning shifts to refining the model’s instruction-following capabilities. We revisit the industrial design application context and, adopting a user-centric approach, reformulate tasks in the paradigm of Visual Question Answering (VQA). Leveraging multimodal language models, we design specific tasks that guide the model to output comprehensive parameter lists and make engineering-critical judgments based on orthographic projections. In this stage, the model is trained to handle the following tasks: 1) Full Parameter List Output Task: Generate a complete and accurate set of parameter key-value pairs— the standard output format for the simulation phase. 2) Parameter Validation Task: Compare extracted parameter key-value pairs with the ground-truth parameter set in the orthographic projections to verify the accuracy of the extracted parameters. 3) View Matching Task: Given two distinct orthographic views, determine if the views correspond to the same 3D object or design based on parameter constraints and rasterized images. 4) Component Counting Task: Identify and count specified components (e.g., bridge piers, columns) in the drawings based on given parameter constraints. This task quantifies design elements to support material estimation and component verification.
Through this post-tuning phase, we strengthen the model’s capacity to interpret and respond to complex design queries, thereby enhancing its interactive reasoning capabilities and increasing its applicability to real-world industrial design workflows.
@@ 5 Experiments
Implementation Details. All experiments were conducted on NVIDIA A100 GPUs. Most experiments used GOT-OCR2.0 as the base model, trained on a single server with 8 A100 GPUs and a batch size of 64. We adopted the AdamW optimizer with a cosine annealing scheduler. The hyperparameters were configured as follows:(1) Learning rate: 1e-6 for RL (GRPO) training and 2e-5 for baseline SFT experiments;(2) Maximum input image size: 2,483,776 pixels;(3) Total GRPO training steps: 1500.
Datasets and Metrics. We evaluated both leading vision–language models (VLMs) and our proposed method on TriView2CAD, following the evaluation strategy outlined in Sec. 3.2. The synthetic test set was used to assess in-domain accuracy. Due to the scarcity of real-world data, all 3,000 real-world samples were reserved exclusively for testing—enabling evaluation of out-of-distribution (OOD) generalization in authentic CAD scenarios.
We design four distinct prompt configurations to probe their orthographic projection reasoning capabilities: 1) Test image‐Only: Models receive only the target rasterized drawings; 2) Test image + Reference Image: In addition to the target drawings, models are given a reference image—rendered on the same geometry layer—with all dimension names to supply semantic alignment; 3) Test image + Answered Pair: Inputs consist of the target drawings plus a correctly paired raster image and its parameter vector, providing an exemplar mapping between visual features and quantitative parameters; 4) Test image + Attribute Explanation: Models are supplied with the target drawings alongside a detailed textual interpretation authored by professional engineers.
Given that orthographic projection reasoning involves complex multi-step inference and current pretrained VLMs have limited exposure to CAD drawings, we incorporated explicit reasoning guidance into each prompt format to encourage coherent reasoning trajectories. For each of the four prompt configurations, we outlined step-by-step instructions for task completion—structured to guide the model through the reasoning process systematically. The complete set of instructions for each format is provided in the supplementary materials.
@@
5.1 Performance on in-domain test set
The experimental results (Tab. 3) demonstrate that our method significantly outperforms existing vision–language models (VLMs) on orthographic projection reasoning tasks across all prompt formats. Specifically, in the "Without Reasoning Guidance" setting, the "+Answered Pair" prompt achieves the highest accuracy of 83.24%. In the "With Reasoning Guidance" setting, our model reaches 84.03% accuracy with the "+Attribute Explanation" prompt. Notably, our method also delivers strong performance under the "Test Image Only" condition, achieving 80.86% accuracy—highlighting the effectiveness of our proposed approach.
@@ 5.1 Performance on in-domain test set
5.2 Performance on real-world test set
We evaluate out-of-distribution (OOD) generalization using all 3,000 real-world samples, which reflect authentic CAD scenarios. The baseline model (Qwen2.5-VL) achieves 13.47% accuracy—lower than its 24.54% performance on synthetic data. This performance degradation is primarily attributed to the more complex structures and occlusion relationships in real-world data (Sec. 3.1), which elevates task difficulty. We also compare with a model trained via standard supervised fine-tuning (SFT), which achieves 36.15% accuracy. This improvement stems from substantial gains in Primitive Counting tasks. However, the model still struggles with Composite Parameter Computation—underscoring the limitations of conventional SFT. Our CReFT-based method achieves 46.67% accuracy, demonstrating its ability to effectively mitigate OOD challenges and deliver substantial performance gains on complex real-world data.
@@ 5.3 Abalation study
With CoT VS without CoT. The results from Tab. 2 reveal a significant improvement in performance when incorporating chain-of-thought (CoT) reasoning. Specifically, the accuracy for individual Task 3 increases from 46.16% to 74.15% after adding CoT , demonstrating a clear performance boost. When combining multiple training tasks, such as Task 1 + Task 3 and Task 2 + Task 3, the improvements are even more pronounced, with accuracies reaching 77.90% and 67.75%, respectively. The most notable enhancement is observed in the combination of Task 1 + Task 2 + Task 3, where the accuracy increases from 46.24% to 81.35%. The primary source of these improvements lies in how CoT helps break down complex Composite Parameter Computation tasks into more manageable, incremental steps. The CoT mechanism, however, guides the model through each individual component of the computation, allowing for a step-by-step approach that improves the model’s reasoning capability.
@@ 6 Conclusion and Limitation
ACKNOWLEDGMENT This work is supported in part by NSFC Project (No. 62176061) and Joint Laboratory of Intelligent Construction Engineering Technology for Operating Railway Lines.
@@ Appendix B Further analysis regarding the three findings
1) Remains a Tough Challenge for pretrained VLMs. Tab. 3 reports performance across four prompt formats without and with reasoning guidance. Overall, orthographic projection reasoning entails not only reading textual dimensions and visual features, but also matching annotations to geometry primitives, counting structural elements, and computing composite parameters. Consequently, none of the off‑the‑shelf models fully masters this composite task. Qwen2.5-VL [46] achieves the highest accuracy—38.88% under reasoning guidance with the Test Image + Attribute Explanation prompt. This improvement stems primarily from its superior ability to parse geometry layers and read dimension labels. Crucially, its performance on reasoning‑intensive parameters remains low. These findings underscore that orthographic projection CAD cannot be solved by prompting alone and requires dedicated fine‑tuning strategy.
@@
3)Introducing reasoning guidance yields consistent gains across all seven VLMs. Introducing reasoning guidance demonstrates that step‐wise reasoning guidance helps models better decompose the multi‐step orthographic projection tasks. In the absence of reasoning guidance, visual exemplar prompts (+Reference Image and +Answered Pair) deliver the largest relative improvements. However, when reasoning guidance is added, the Attribute Explanation prompt shows the greatest uplift. It stems from reasoning guidance is textual form, which synergizes most effectively with textual inputs. Hence, models with stronger text encoders (e.g., DeepSeek‑V1) exhibit disproportionately larger boosts. These results underscore the necessity of integrating structured reasoning guidance to advance orthographic projection CAD reasoning.